OpenStudy (anonymous):

the perimeter of the rectangle equals 372. Find the value of x

OpenStudy (anonymous):

well find two number that you can divide by two that add up to 372

OpenStudy (anonymous):

the numbers of the rectangle is 3x-7 and 5x+9

OpenStudy (anonymous):

is one the width and the other the length

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6x-14+10x+18= 372 16x + 4 = 372 16x = 368 x=23

OpenStudy (anonymous):

now just plug 23 into the original equations 3(23)+7 69+7 76 5(23)+9 115+9 124

OpenStudy (anonymous):

ok thanks. I have a another question. The diameter of a cylinder is 7yd. The height is 4yd. Find the volume of the cylinder in terms of pie?

OpenStudy (raden):

Volume = pi * r^2 * h (r = radius, h = height)

OpenStudy (anonymous):

can you help me with the answer?

OpenStudy (raden):

Volume = pi * r^2 * h Volume = pi * (3.5)^2 * 4 = ... pi notice that r = diameter/2

OpenStudy (anonymous):

can you just give me the answer please?

OpenStudy (raden):

just calculate this 3.5 x 3.5 x 4 = ... ? (IF You not sure, use the calculator)